# From Consistency to Complementarity: Aligned and Disentangled Multi-modal Learning for Time Series Understanding and Reasoning

### v1

@@ Abstract
Advances in multi-modal large language models (MLLMs) have inspired time series understanding and reasoning tasks, that enable natural language querying over time series, producing textual analyses of complex temporal dynamics. Recent attempts hybridize numerical time series with their visualized plots, facilitating precise value reasoning and visual structure comprehension for comprehensive time series understanding of MLLMs. However, effective cross-modal integration remains challenging due to fine-grained temporal misalignment across modalities and severe entanglement between shared and modality-specific semantics, which hinder localized interpretation and complementary reasoning. To address these issues, we propose MADI, a multi-modal LLM enhanced with fine-grained alignment and disentangled interaction, featuring (1) Patch-level Alignment, which enforces physically grounded fine-grained correspondence across heterogeneous modalities, (2) Discrete Disentangled Interaction, which separates modality-common semantics into compact discrete latents and adaptively synergizes the purified modality-unique information, and (3) Critical-token Highlighting, which emphasizes informative, query-relevant signals for robust reasoning. Experiments on synthetic and real-world benchmarks show that MADI consistently outperforms general-purpose LLMs and time-series-specialized MLLMs.
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), TSUR tasks have become dominated by LLM-based paradigms, enabling flexible user queries and textual outputs. Beyond linguistic inputs, existing works primarily focus on how to formulate time series into effective observational modalities that can be jointly processed with language by LLMs. Numerical-centric methods directly feed raw time series values into LLMs as an additional modality (Wang et al., 2025a, b; Xie et al., 2025). While preserving precision, studies indicate that LLMs struggle to recognize high-level temporal structures like trends and periodicity (Liu et al., 2025c), reflecting a fundamental numerical-language modality gap induced by text-only LLM pre-training (Tan et al., 2024). In contrast, visual-centric methods convert time series into visual plots and leverage the strong visual processing abilities of multi-modal LLMs (MLLMs) (Liu et al., 2025c; Merrill et al., 2024; Zhang et al., 2025a). Although effective at capturing high-level temporal structures, these approaches often sacrifice fine-grained numerical details that are critical for precise temporal analysis (Xu et al., 2025; Zhao et al., 2025). Recent hybrid approaches like GEM (Lan et al., 2025) attempt to combine numerical and visual modalities within MLLMs via token concatenation for comprehensive time series understanding. Despite these advances, effectively integrating these two observational modalities for MLLMs reveals two critical challenges:
@@ 2.2 Time Series Understanding and Reasoning
Beyond classical analysis, time series understanding and reasoning (TSUR) emphasizes interpretable analyses of temporal dynamics under flexible natural language queries, leveraging LLMs. Existing approaches can be broadly classified as numerical-centric, visual-centric, and hybrid paradigms. Numerical-centric methods represent time series values as sequences of textual or continuous tokens (Wang et al., 2025a, b; Xie et al., 2025), preserving precision but often struggling with high-level temporal structures due to the text-centric pre-training of LLMs (Tan et al., 2024; Liu et al., 2025c). Visual-centric approaches (Liu et al., 2025c; Merrill et al., 2024; Zhang et al., 2025a) reframe time series analysis as a visual question-answering task, transforming the series into plots and leveraging advanced MLLMs (Liu et al., 2025c; Merrill et al., 2024; Zhang et al., 2025a). These methods can effectively capture high-level patterns, e.g., trend and periodicity, but sacrifice fine-grained numerical details (Xu et al., 2025; Zhao et al., 2025). Hybrid approaches, e.g., GEM (Lan et al., 2025), combine both modalities, but typically rely on simple concatenation and remain domain-specific (e.g., ECG), limiting the principled cross-modal alignment and interaction. Our work addresses these limitations with a generalizable MLLM that integrates numerical precision and visual abstractions via fine-grained alignment and disentangled cross-modal interaction, supporting diverse domains and flexible query types.
@@ 4 Methodology
Figure 2 illustrates the overall architecture of MADI. It is built upon a pre-trained MLLM, reusing its vision encoder and LLM to perform time series understanding and reasoning (TSUR). Specifically, Patch-level Alignment (PA) first expands each time series as patch-aligned numerical, visual and textual modalities, aligning them in a consistent embedding space. Based on these aligned tokens, Discrete Disentangled Interaction (DDI) separates modality-shared and unique semantics for numerical and visual inputs and synergizes their complementarities. Critical-token Highlighting (CTH) then identifies informative, question-relevant tokens from the PA-aligned embeddings and prepends them to the DDI-fused multi-modal sequence. The resulting multi-modal tokens are concatenated with the textual input tokens and decoded by the LLM to generate the output.
@@ 4.1.3 Contrastive Alignment
where sim denotes cosine similarity, sg denotes the stop-gradient operator, $\tau$ is a temperature parameter, and $(i,j^{\prime})$ indexes all visual tokens within the $i$-th time series instance. The numerical–caption alignment loss $\mathcal{L}^{n\text{-}s}_{\text{align}}$ is computed analogously. We stop the gradients of encoded visual and textual tokens to optimize the numerical encoder only. The overall alignment objective is: $\mathcal{L}_{\text{PA}}=\mathcal{L}^{n\text{-}v}_{\text{align}}+\mathcal{L}^{n\text{-}s}_{\text{align}}$.
@@ 4.2.1 Discrete Modality Disentanglement
A straightforward approach to identify modality-unique semantics is to apply conventional disentanglement techniques (Qian et al., 2025), which use separate projection heads to extract modality-common and modality-specific representations, alongside regularization encouraging high correlation among common components and low correlation with unique components. However, existing methods (Xia et al., 2023; Huang et al., 2025) operate in continuous embedding spaces, which are unbounded and prone to capturing redundant or noisy correlations. In contrast, we perform disentanglement in a discrete latent space using vector quantization (VQ), which enforces compactness and consistency for modality-common semantics. Given numerical and visual tokens $\mathbf{E}_{i}^{n},\mathbf{E}_{i}^{v}\in\mathbb{R}^{\tilde{T}_{i}\times D}$, we employ a shared hierarchical discretization block to extract modality-common representations $\mathbf{Z}_{i}^{n}$ and $\mathbf{Z}_{i}^{v}$. The modality-unique components are then obtained by residual decomposition:
@@
Hierarchical Discretization. To capture the multi-scale time series semantics, we propose a hierarchical RVQ strategy featured by $M$ hierarchical codebooks $\{\mathcal{C}^{(m)}\}_{m=1}^{M}$, where the $m$-th codebook $\mathcal{C}^{(m)}$ contains $K\cdot 2^{m-1}$ codes and operates at a temporal resolution downsampled by $2^{M-m}$ tokens. Specifically, we first follow (Yu et al., 2022) and project continuous tokens $\mathbf{E}_{i}^{n}\in\mathbb{R}^{\tilde{T}_{i}\times D}$ to a lower-dimensional space via $\tilde{\mathbf{E}}_{i}^{n}=\text{Linear}(\mathbf{E}_{i}^{n})\in\mathbb{R}^{\tilde{T}_{i}\times d}$, which increases the density of codes and reduces the quantization error. At each quantization stage $m$, given the residual token sequence $\mathbf{R}_{i}^{n;(m-1)}=[\mathbf{r}_{i;1}^{n;(m-1)},\dots,\mathbf{r}_{i;\tilde{T}_{i}}^{n;(m-1)}]^{\top}\in\mathbb{R}^{\tilde{T}_{i}\times d}$ from the last round, we partition it into non-overlapping segments of length $W=2^{M-m}$, and compute segment-level representations by average pooling over the tokens within each segment. These pooled representations are then broadcasted back to all token positions in the corresponding segment. This process can be formulated as $\tilde{\mathbf{r}}_{i;j}^{n;(m-1)}=\frac{1}{W}\sum_{b=j^{*}}^{j^{*}+W}\mathbf{r}_{i;b}^{n;(m-1)}$, where $j^{*}=[(j-1)/W]$ denotes the segment index of the $j$-th token. Afterwards, we follow the steps illustrated in Section 3.2, iteratively quantizing the residual features, from coarse to fine scales:
@@
where the initial residual tokens are $\mathbf{R}_{i}^{n;(0)}=\tilde{\mathbf{E}}_{i}^{n}$. The final discretized features are accumulated by $\mathbf{q}_{i;j}^{n}=\sum_{m=1}^{M}\mathbf{q}_{i;j}^{n;(m)}$. After quantization, the discrete representations should be projected back to $D$-dimensional space via $\mathbf{z}_{i;j}^{n}=\text{Linear}(\mathbf{q}_{i;j}^{n})\in\mathbb{R}^{D}$, which serve as the modality-common tokens, and constitute $\mathbf{Z}_{i}^{n}\in\mathbb{R}^{\tilde{T}_{i}\times d}$. The modality-common tokens of visual time series $\mathbf{Z}_{i}^{v}$ are derived in a similar way. The qunatization process is supervised the $\mathcal{L}_{\text{vq}}$ illustrated in Section 3.2.
Disentanglement Regularization. To effectively decouple modality-common and modality-unique information, we introduce two regularization terms. First, to align modality-common semantics, we apply a bidirectional contrastive objective between quantized numerical and visual tokens: $\mathcal{L}_{\text{com}}=\frac{1}{2}(\mathcal{L}^{n\text{-}v}_{\text{com}}+\mathcal{L}^{v\text{-}n}_{\text{com}})$, where
@@
To suppress leakage of shared semantics into modality-unique components, we enforce orthogonality between the common and unique tokens:
@@ 4.3 Critical-token Highlighting
Although PA and DDI yield aligned and complementary representations, directly feeding them into the LLM may dilute question-relevant or inherently important information, limiting generalization. We propose a Critical-token Highlighting (CTH) module, which retains all original tokens and prepends several informative tokens to explicitly emphasize key signals, conditioned on both the question-aware and modality-intrinsic saliency. Specifically, the question $Q$ is first tokenized and embedded as $\mathbf{E}^{q}=\text{LLM-Embed}(Q)\in\mathbb{R}^{N^{q}\times D}$. Given question embeddings $\mathbf{E}^{q}$, and continuous numerical/visual tokens $\mathbf{E}_{i}^{o}\in\mathbb{R}^{\tilde{T}_{i}\times D}$ ($o\in\{n,v\}$), CTH employs two parallel cross-attention branches.
Question-conditioned Branch. We first compress the variable-length question tokens into $H$ learnable query tokens $\mathbf{Q}^{q}$ via an attention-based pooling layer, and then use the compressed question-aware queries to attend over the modality tokens, which can be formulated as:
@@
The highlighted tokens are fused as $\mathbf{H}_{i}^{o}=\mathbf{H}^{o;q}_{i}+\mathbf{H}^{o;s}_{i}$ and prepended to the modality token sequence via $\hat{\mathbf{E}}^{o}_{i}=\text{Concat}(\mathbf{H}_{i}^{o},\overline{\mathbf{E}}^{o}_{i})\in\mathbb{R}^{(H+\tilde{T}_{i})\times D},\ o\in\{n,v\}$. This results in holistic and question-aware modality representations for LLM understanding and reasoning.
@@ 4.4 LLM Backbone and Training Objective
The enhanced numerical and visual tokens $\{\hat{\mathbf{E}}^{n}_{i},\hat{\mathbf{E}}^{v}_{i}\}_{i=1}^{|\mathcal{X}|}$ are concatenated with textual input tokens, including the question tokens $\mathbf{E}^{q}$ and context tokens $\mathbf{E}^{c}$ (derived from the context $C$ using the LLM’s embeddings), via a position-aware concatenation operator (Xie et al., 2025):
@@
where $\text{PosConcat}(\cdot)$ integrates multi-modal tokens with the question tokens according to the positions of the time series instances in the original input. The fused sequence $\mathbf{E}^{*}$ is fed into the backbone LLM to generate the answer $A$ autoregressively: $P(A\mid\mathbf{E}^{*})=\prod_{b=1}^{|A|}P(A_{b}\mid A_{<b},\mathbf{E}^{*})$, where $A_{b}$ denotes the $b$-th token of the target answer sequence. In this study, we adopt Qwen2.5-VL-7B-Instruct (Bai et al., 2025) as our backbone MLLM. For model training, the language modeling loss is: $\mathcal{L}_{\text{LM}}=-\sum_{b=1}^{|A|}\log P(A_{b}\mid A_{<b},\mathbf{E}^{*})$, and the overall training objective incorporates the regularization terms from PA and DDI: $\mathcal{L}=\mathcal{L}_{\text{LM}}+\lambda_{1}\cdot\mathcal{L}_{\text{PA}}+\lambda_{2}\cdot\mathcal{L}_{\text{DDI}}$, where $\lambda_{1},\lambda_{2}$ denote the loss weights.

### v2

@@ Abstract
Advances in multi-modal large language models (MLLMs) have inspired time series understanding and reasoning tasks, that enable natural language querying over time series, producing textual analyses of complex temporal dynamics. Recent attempts hybridize numerical time series with their visualized plots, facilitating precise value reasoning and visual structure comprehension for comprehensive time series understanding of MLLMs. However, effective numerical-visual modality integration remains challenging due to fine-grained temporal misalignment across modalities and severe entanglement between shared and modality-specific semantics, which hinder localized interpretation and complementary reasoning. To address these issues, we propose MADI, a multi-modal LLM enhanced with fine-grained alignment and disentangled interaction, featuring (1) Patch-level Alignment, which enforces physically grounded fine-grained correspondence across heterogeneous modalities, (2) Discrete Disentangled Interaction, which separates modality-common semantics into compact discrete latents and adaptively synergizes the purified modality-unique information, and (3) Critical-token Highlighting, which emphasizes informative, query-relevant signals for robust reasoning. Experiments on synthetic and real-world benchmarks show that MADI consistently outperforms general-purpose LLMs and time-series-specialized MLLMs.
@@ 1 Introduction
With the rapid advancement of large language models (LLMs), TSUR tasks have become dominated by LLM-based paradigms, enabling flexible user queries and textual outputs. Beyond linguistic inputs, existing works primarily focus on how to formulate time series into effective observational modalities that can be jointly processed with language by LLMs. Numerical-centric methods directly feed raw time series values into LLMs, either by encoding them as plain text via the LLM tokenizer and embeddings (Liu et al., 2025c; Merrill et al., 2024), or by introducing specialized time series encoders as an additional modality (Wang et al., 2025a, b; Xie et al., 2025). While preserving precision, studies indicate that LLMs struggle to recognize high-level temporal structures like trends and periodicity (Liu et al., 2025c), reflecting a fundamental numerical-language modality gap induced by text-only LLM pre-training (Tan et al., 2024). In contrast, visual-centric methods convert time series into visual plots and leverage the strong visual processing abilities of multi-modal LLMs (MLLMs) (Liu et al., 2025c; Merrill et al., 2024; Zhang et al., 2025a). Although effective at capturing high-level temporal structures, these approaches often sacrifice fine-grained numerical details that are critical for precise temporal analysis (Xu et al., 2025; Zhao et al., 2025). Recent hybrid approaches like GEM (Lan et al., 2025) attempt to combine numerical and visual modalities within MLLMs via token concatenation for comprehensive time series understanding. Despite these advances, effectively integrating these two observational modalities for MLLMs reveals two critical challenges:
@@ 2.2 Time Series Understanding and Reasoning
Beyond classical analysis, time series understanding and reasoning (TSUR) emphasizes interpretable analyses of temporal dynamics under flexible natural language queries, leveraging LLMs. Existing approaches can be broadly classified as numerical-centric, visual-centric, and hybrid paradigms. Numerical-centric methods represent time series values as sequences of textual or time-series-specific tokens (Merrill et al., 2024; Wang et al., 2025a, b; Xie et al., 2025), preserving precision but often struggling with high-level temporal structures due to the text-centric pre-training of LLMs (Tan et al., 2024; Liu et al., 2025c). Visual-centric approaches (Liu et al., 2025c; Merrill et al., 2024; Zhang et al., 2025a) reframe time series analysis as a visual question-answering task, transforming the series into plots and leveraging advanced MLLMs (Liu et al., 2025c; Merrill et al., 2024; Zhang et al., 2025a). These methods can effectively capture high-level patterns, e.g., trend and periodicity, but sacrifice fine-grained numerical details (Xu et al., 2025; Zhao et al., 2025). Hybrid approaches, e.g., GEM (Lan et al., 2025), combine both modalities, but typically rely on simple concatenation and remain domain-specific (e.g., ECG), limiting the principled cross-modal alignment and interaction. Our work addresses these limitations with a generalizable MLLM that integrates numerical precision and visual abstractions via fine-grained alignment and disentangled cross-modal interaction, supporting diverse domains and flexible query types.
@@ 4 Methodology
Figure 2 illustrates the overall architecture of MADI. It is built upon a pre-trained MLLM, reusing its vision encoder and LLM to perform time series understanding and reasoning (TSUR). Specifically, Patch-level Alignment (PA) first expands each time series as patch-aligned numerical, visual and textual modalities, aligning them in a consistent embedding space. Based on these aligned token embeddings, Discrete Disentangled Interaction (DDI) separates modality-shared and unique semantics for numerical and visual inputs and synergizes their complementarities. Critical-token Highlighting (CTH) then identifies informative, question-relevant tokens from the PA-aligned embeddings and prepends them to the DDI-fused multi-modal sequence. The resulting multi-modal embeddings are concatenated with the textual input embeddings and decoded by the LLM to generate the output.
@@ 4.1.3 Contrastive Alignment
where sim denotes cosine similarity, sg denotes the stop-gradient operator, $\tau$ is a temperature parameter, and $(i,j^{\prime})$ indexes all visual tokens within the $i$-th time series instance. The numerical–caption alignment loss $\mathcal{L}^{n\text{-}s}_{\text{align}}$ is computed analogously. We stop the gradients of encoded visual and textual embeddings to optimize the numerical encoder only. The overall alignment objective is: $\mathcal{L}_{\text{PA}}=\mathcal{L}^{n\text{-}v}_{\text{align}}+\mathcal{L}^{n\text{-}s}_{\text{align}}$.
@@ 4.2.1 Discrete Modality Disentanglement
A straightforward approach to identify modality-unique semantics is to apply conventional disentanglement techniques (Qian et al., 2025), which use separate projection heads to extract modality-common and modality-specific representations, alongside regularization encouraging high correlation among common components and low correlation with unique components. However, existing methods (Xia et al., 2023; Huang et al., 2025) operate in continuous embedding spaces, which are unbounded and prone to capturing redundant or noisy correlations. In contrast, we perform disentanglement in a discrete latent space using vector quantization (VQ), which enforces compactness and consistency for modality-common semantics. Given numerical and visual token embeddings $\mathbf{E}_{i}^{n},\mathbf{E}_{i}^{v}\in\mathbb{R}^{\tilde{T}_{i}\times D}$, we employ a shared hierarchical discretization block to extract modality-common representations $\mathbf{Z}_{i}^{n}$ and $\mathbf{Z}_{i}^{v}$. The modality-unique components are then obtained by residual decomposition:
@@
Hierarchical Discretization. To capture the multi-scale time series semantics, we propose a hierarchical RVQ strategy featured by $M$ hierarchical codebooks $\{\mathcal{C}^{(m)}\}_{m=1}^{M}$, where the $m$-th codebook $\mathcal{C}^{(m)}$ contains $K\cdot 2^{m-1}$ codes and operates at a temporal resolution downsampled by $2^{M-m}$ tokens. Specifically, we first follow (Yu et al., 2022) and project continuous embeddings $\mathbf{E}_{i}^{n}\in\mathbb{R}^{\tilde{T}_{i}\times D}$ to a lower-dimensional space via $\tilde{\mathbf{E}}_{i}^{n}=\text{Linear}(\mathbf{E}_{i}^{n})\in\mathbb{R}^{\tilde{T}_{i}\times d}$, which increases the density of codes and reduces the quantization error. At each quantization stage $m$, given the residual token sequence $\mathbf{R}_{i}^{n;(m-1)}=[\mathbf{r}_{i;1}^{n;(m-1)},\dots,\mathbf{r}_{i;\tilde{T}_{i}}^{n;(m-1)}]^{\top}\in\mathbb{R}^{\tilde{T}_{i}\times d}$ from the last round, we partition it into non-overlapping segments of length $W=2^{M-m}$, and compute segment-level representations by average pooling over the tokens within each segment. These pooled representations are then broadcasted back to all token positions in the corresponding segment. This process can be formulated as $\tilde{\mathbf{r}}_{i;j}^{n;(m-1)}=\frac{1}{W}\sum_{b=j^{*}}^{j^{*}+W}\mathbf{r}_{i;b}^{n;(m-1)}$, where $j^{*}=[(j-1)/W]$ denotes the segment index of the $j$-th token. Afterwards, we follow the steps illustrated in Section 3.2, iteratively quantizing the residual features, from coarse to fine scales:
@@
where the initial residual token embeddings are $\mathbf{R}_{i}^{n;(0)}=\tilde{\mathbf{E}}_{i}^{n}$. The final discretized features are accumulated by $\mathbf{q}_{i;j}^{n}=\sum_{m=1}^{M}\mathbf{q}_{i;j}^{n;(m)}$. After quantization, the discrete representations should be projected back to $D$-dimensional space via $\mathbf{z}_{i;j}^{n}=\text{Linear}(\mathbf{q}_{i;j}^{n})\in\mathbb{R}^{D}$, which serve as the modality-common tokens, and constitute $\mathbf{Z}_{i}^{n}\in\mathbb{R}^{\tilde{T}_{i}\times d}$. The modality-common token embeddings of visual time series $\mathbf{Z}_{i}^{v}$ are derived in a similar way. The qunatization process is supervised the $\mathcal{L}_{\text{vq}}$ illustrated in Section 3.2.
Disentanglement Regularization. To effectively decouple modality-common and modality-unique information, we introduce two regularization terms. First, to align modality-common semantics, we apply a bidirectional contrastive objective between quantized numerical and visual token embeddings: $\mathcal{L}_{\text{com}}=\frac{1}{2}(\mathcal{L}^{n\text{-}v}_{\text{com}}+\mathcal{L}^{v\text{-}n}_{\text{com}})$, where
@@
To suppress leakage of shared semantics into modality-unique components, we enforce orthogonality between the common and unique token embeddings:
@@ 4.3 Critical-token Highlighting
Although PA and DDI yield aligned and complementary representations, directly feeding them into the LLM may dilute question-relevant or inherently important information, limiting generalization. We propose a Critical-token Highlighting (CTH) module, which retains all original tokens and prepends several informative tokens to explicitly emphasize key signals, conditioned on both the question-aware and modality-intrinsic saliency. Specifically, the question $Q$ is first tokenized and embedded as $\mathbf{E}^{q}=\text{LLM-Embed}(Q)\in\mathbb{R}^{N^{q}\times D}$. Given question embeddings $\mathbf{E}^{q}$, and continuous numerical/visual embeddings $\mathbf{E}_{i}^{o}\in\mathbb{R}^{\tilde{T}_{i}\times D}$ ($o\in\{n,v\}$), CTH employs two parallel cross-attention branches.
Question-conditioned Branch. We first compress the variable-length question embeddings into $H$ learnable query embeddings $\mathbf{Q}^{q}$ via an attention-based pooling layer, and then use the compressed question-aware queries to attend over the modality tokens, which can be formulated as:
@@
The highlighted token embeddings are fused as $\mathbf{H}_{i}^{o}=\mathbf{H}^{o;q}_{i}+\mathbf{H}^{o;s}_{i}$ and prepended to the modality token sequence via $\hat{\mathbf{E}}^{o}_{i}=\text{Concat}(\mathbf{H}_{i}^{o},\overline{\mathbf{E}}^{o}_{i})\in\mathbb{R}^{(H+\tilde{T}_{i})\times D},\ o\in\{n,v\}$. This results in holistic and question-aware modality representations for LLM understanding and reasoning.
@@ 4.4 LLM Backbone and Training Objective
The enhanced numerical and visual token embeddings $\{\hat{\mathbf{E}}^{n}_{i},\hat{\mathbf{E}}^{v}_{i}\}_{i=1}^{|\mathcal{X}|}$ are concatenated with textual input tokens, including the question embeddings $\mathbf{E}^{q}$ and context embeddings $\mathbf{E}^{c}$ (derived from the context $C$ using the LLM’s embeddings), via a position-aware concatenation operator (Xie et al., 2025):
@@
where $\text{PosConcat}(\cdot)$ integrates multi-modal token embeddings in the textual inputs according to the positions of the time series instances in the original input. The fused sequence $\mathbf{E}^{*}$ is fed into the backbone LLM to generate the answer $A$ autoregressively: $P(A\mid\mathbf{E}^{*})=\prod_{b=1}^{|A|}P(A_{b}\mid A_{<b},\mathbf{E}^{*})$, where $A_{b}$ denotes the $b$-th token of the target answer sequence. In this study, we adopt Qwen2.5-VL-7B-Instruct (Bai et al., 2025) as our backbone MLLM. For model training, the language modeling loss is: $\mathcal{L}_{\text{LM}}=-\sum_{b=1}^{|A|}\log P(A_{b}\mid A_{<b},\mathbf{E}^{*})$, and the overall training objective incorporates the regularization terms from PA and DDI: $\mathcal{L}=\mathcal{L}_{\text{LM}}+\lambda_{1}\cdot\mathcal{L}_{\text{PA}}+\lambda_{2}\cdot\mathcal{L}_{\text{DDI}}$, where $\lambda_{1},\lambda_{2}$ denote the loss weights.